# Ordinary primes for $\operatorname{GL}_{2}$-type abelian varieties and weight $2$ modular forms

### v1

@@ 1. Introduction
Let $A$ be an abelian variety of dimension $g$ over a number field $F$. Let $v$ be a prime of $F$ that is also a prime of good reduction for $A$. We denote by $A_{v}$ the reduction of $A$ at $v$ and by $P_{A,v}(X)$ the characteristic polynomial of the Frobenius endomorphism of $A_{v}$. Then, $P_{A,v}(X)$ is a polynomial of degree $2g$ with integer coefficients. We write the coefficient of the middle term $X^{g}$ in $P_{A,v}(X)$ as $a_{v,g}(A)$. Then, the prime $v$ is an ordinary prime of $A$ over $F$ if $p\nmid a_{v,g}(A)$, where $p$ is the rational prime below $v$. In this case, $A_{v}$ is an ordinary abelian variety over $\mathbb{F}_{v}$ and we say $A$ is ordinary at $v$. For example, if $A$ is an elliptic curve over $\mathbb{Q}$, then a rational prime $v=p$ is an ordinary prime of $A$ if and only if $p\nmid a_{p,1}(A)$, where $a_{p,1}(A)=p+1-|A_{p}(\mathbb{F}_{p})|$ is the Frobenius trace of $A$ at $p$. When $p\geq 5$, this is also equivalent to that $a_{p,1}(A)\neq 0$ by the Hasse–Weil bound.
@@
A comparable definition of “ordinary” also appears in the study of modular forms. Let $f$ be a newform of weight $k\geq 2$ and level $N$, let $a_{n}(f)$ be the $n$-th Fourier coefficient of $f$, and let $K_{f}$ be the coefficient field of $f$. For a rational prime $p\nmid N$ and a prime $\lambda|p$ of $K_{f}$, $f$ is called $\lambda$-ordinary if $\lambda\nmid(a_{p}(f))$, and $f$ is called $p$-ordinary if $f$ is $\lambda$-ordinary for all $\lambda|p$, or equivalently, $\lambda\nmid(a_{p}(f))$ for all primes $\lambda|p$ of $K_{f}$. We will use the phrases “$f$ is $p$-ordinary”, “$f$ is ordinary at $p$”, and “$p$ is an ordinary prime of $f$ over $\mathbb{Q}$” interchangeably.
The notion of ordinary modular forms first appeared in a paper by Hida [Hid81]. This paper was followed by a series work of Hida, where he established the theory of ordinary $p$-adic modular forms, or nowadays known as Hida theory. The ordinary assumption has since played an important role in the theory of ($p$-adic) modular forms, particularly in Iwasawa theory, e.g., in Emerton–Pollack–Weston [EPW06] and Skinner–Urban [SU14] (and many others). The definition of ordinariness that we adopt in this paper is, however, different from (but closely related to) Hida’s definition. Roughly speaking, Hida’s definition of $p$-ordinariness only requires $a_{p}(f)$ to be a $p$-adic unit under a prechosen embedding $K_{f}\hookrightarrow\overline{\mathbb{Q}_{p}}$, while our definition requires $a_{p}(f)$ to be a $p$-adic unit under all embeddings $K_{f}\hookrightarrow\overline{\mathbb{Q}_{p}}$. In particular, our definition of $p$-ordinariness coincides with the one by Hida if $K_{f}=\mathbb{Q}$. The reason for us to adopt a different definition is that a weight $2$ newform $f$ is ordinary at $p$ under our definition if and only if the abelian variety $A_{f}/\mathbb{Q}$ associated to $f$ via the Eichler–Shimura theory is ordinary at $p$ (see Proposition 3.12).
In this paper, we prove a positive density result for the sets of ordinary primes of certain abelian varieties and modular forms by analyzing under the framework of Galois representations.
We begin by reviewing prior works on the study of ordinary primes for an abelian variety $A$ over a number field $F$. A conjecture commonly attributed to Serre [DMOS82, Proposition 2.6, p. 370] predicts that there are infinitely many ordinary primes for $A/F$. Furthermore, it is conjectured that the density of ordinary primes for $A$ over $\overline{F}$ is 1. The conjecture is proved when the dimension of $A$ is 1 or 2 [Ser00, DMOS82, Saw16]. However, much less is known for higher dimensional abelian varieties. Nonetheless, the conjecture holds for special cases where $A$ is not generic, such as the case when $A/F$ is a CM abelian variety [Fit24, Remark 12] (i.e., the geometric endomorphism algebra $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ is a CM field of degree $2g$) or when the Mumford–Tate group of $A$ is small in the sense of [Noo95, Theorem 2.2, p. 168] or [Pin98, Theorem 7.1].
A recent work by Fité [Fit24, Theorem 3] showed that the sets of ordinary primes of certain modular abelian varieties have positive density. More generally, Fité [Fit24, Theorem 1, Theorem 4, and Corollary 5] proved this result for abelian varieties that have additional endomorphisms. In comparison with Fité’s work on modular abelian varieties, the work by Suh [Suh20, Theorem, p. 649] showed that the sets of ordinary primes of certain Hilbert newforms of parallel weight 2 have positive density. The reason for this comparison is that Hilbert newforms of parallel weight $2$ are conjecturally associated to certain abelian varieties over totally real number fields with extra endomorphisms. More recently, the forthcoming work by Cantoral Farfán, Li, Mantovan, Pries, and Tang established a positive density result for the sets of ordinary primes of certain abelian varieties parametrized by unitary Shimura curves with simple signature.
We extend the family of higher dimensional abelian varieties whose sets of ordinary primes have positive density. Our first two results are for $\operatorname{GL}_{2}$-type abelian varieties over $\mathbb{Q}$ and quadratic number fields of specific dimensions.
@@
More generally, let $A/F$ be a simple abelian variety of an arbitrary dimension $g$ with whose endomorphism algebra contains a number field $K$ of degree $g$. The second result gives an affirmative answer to the positive density conjecture on ordinary primes for $A$ over $F$ under certain restrictions on $F$ and $K$.
@@
    $A$ is absolutely simple, i.e., $A/\overline{F}$ is simple;
@@
    $A/\overline{F}$ is of the first kind, i.e., $A/\overline{F}$ is not of type IV in the Albert’s classification.
@@
The family we consider is taken from [TTV91]. Let $p$ be a rational prime with $p\neq 2$ or $5$. Let $\zeta_{p}$ denote the primitive $p$-th root of unity in $\overline{\mathbb{Q}}$, $g(X)\in\mathbb{Z}[X]$ be the minimal polynomial of $-(\zeta_{p}+\zeta_{p}^{-1})$, and $f_{t}(X):=Xg(X^{2}-2)+t\in\mathbb{Z}[X]$ for any $t\in\overline{\mathbb{Q}}$. The equation
@@
characterizes a family of absolutely simple hyperelliptic curves of genus $\frac{p-1}{2}$ [TTV91, Corollary 6]. For each curve $C_{t}$, the Jacobian variety $J_{C_{t}}$ is defined over $F:=\mathbb{Q}(t)$, of dimension $\frac{p-1}{2}$, and contains the totally real field $K:=\mathbb{Q}(\zeta_{p}+\zeta_{p}^{-1})$ in its endomorphism algebra [TTV91, Theorem 1]. Note that $[K:\mathbb{Q}]=\frac{p-1}{2}=\dim J_{C_{t}}$. Since the Galois extension $K/\mathbb{Q}$ is cyclic when $p\geq 3$, if $\ell$ is inert in $K$, then the order of $\ell\pmod{p}$ is divisible by $\frac{p-1}{2}$, and if $\ell$ splits into two primes of the same inertia degree, then the order of $\ell\pmod{p}$ is divisible by $\frac{p-1}{4}$ (if $p\equiv 1\pmod{4}$). Clearly, such a prime $\ell$ exists. Thus, the condition (P1) is satisfied if $p\geq 3$, and the condition (P2) is satisfied if $p\equiv 1\pmod{4}$. Finally, by choosing a number field $F$ such that $\ell$ lies below a degree $1$ prime of $F$ and taking $t$ as the primitive element of $F$, we see that $\{J_{C_{t}}\}_{t\in F}$ is a family of abelian varieties whose sets of ordinary primes over $F$ have positive density.
In the case of modular forms, we obtain the following result under our definition of ordinary primes of modular forms.
@@
Let $f$ be a newform of weight $2$ and level $N$ and let $K_{f}$ denote the coefficient field of $f$. Suppose that $[K_{f}:\mathbb{Q}]=q$ or $2q$ for some prime $q$. Then, the set of ordinary primes of $f$ over $\mathbb{Q}$ has positive density.
@@
To prove Theorem 1.1, Theorem 1.5 in the case of “$2q$” and Theorem 1.3 in the case of (P2), we begin by proving Theorem 1.6, a general density result for compatible systems of Galois representations. We then apply this theorem to the $\ell$-adic Galois representations of abelian varieties and modular forms.
@@
    (Ramanujan’s conjecture/Hasse–Weil bound) the sequence $(a_{v})_{v\in\Sigma_{F}\setminus S}$ satisfies that
@@
The general approach to proving such results can be traced back to the work of Serre, Katz, and Ogus, and has been successfully adapted with modifications by Pink, Noot, and others, as previously mentioned. Our proof of Theorem 1.6 is also similar and we sketch it as follows. First, we demonstrate that there is a positive density set $S_{0}$ of degree 1 primes $v$ of $F$, lying above rational primes $p$, that split into two distinct primes in $K$ of the same inertia degree. This requires analysis of the splitting behaviors of primes in extensions of number fields (that are not necessarily Galois). Next, we associate to each prime $v\in S_{0}$ the value $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}$ and prove that if $v\notin S^{\mathrm{ord}}$, then $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}$ belongs to a finite set (with size at most $12^{2g}+$ 1). The proof of finiteness relies on the splitting behavior of $p$ and the bound bound on the archimedean norm of $a_{v}$. Finally, we perform explicit computations using the images of residual Galois representations and the Chebotarev density theorem to show that the density of the primes $v\in S_{0}$ with $v\notin S^{\mathrm{ord}}$ is 0. This step involves studying the images of the $\lambda$-adic Galois representations of abelian varieties and modular forms.
Though Theorem 1.1, Theorem 1.3, and Theorem 1.5 are all applications of Theorem 1.6, we will present a simpler proof of Theorem 1.1 by using the modularity of $\operatorname{GL}_{2}$-type abelian varieties over $\mathbb{Q}$ and Theorem 1.5. Additionally, we emphasize that the surjectivity condition (4) in Theorem 1.6 is intentionally relaxed to avoid the need for assuming that the endomorphisms of $A$ are defined over the base field and that the modular forms have no inner twists.
On the other hand, the proofs of Theorem 1.1, Theorem 1.2, and Theorem 1.5 in the case of “$q$”, and the proof of Theorem 1.3 in the case of (P1) are much simpler. In this case, we only need the following density result for a sequence of algebraic integers indexed by primes.
@@
    (Ramanujan’s conjecture) $|\iota(a_{v})|_{\mathbb{C}}\leq 2\mathrm{N}(v)^{1/2}$ for all $v\in\Sigma_{F}$ with $v\notin S$ and all embeddings $\iota:K\hookrightarrow\mathbb{C}$;
@@
    (non-lacunary) the set $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$ of primes of $F$ has density 00.
@@
One can also apply Theorem 1.6 and Theorem 1.7 to obtain similar positive density results on the ordinary primes of certain Hilbert newforms, where one assumes a condition similar to that in Theorem 1.3 on the existence of some rational prime with specific joint splitting behaviors in the base field and in the coefficient field of the Hilbert newform. The main ingredients for applying Theorem 1.6 and Theorem 1.7 to the case of Hilbert modular forms are given by [Car86] and [Tay89] on the existence of the associated Galois representations, [Nek12, Appendix B] on the surjectivity of the Galois images, and [Bla06] on Ramanujan’s conjecture.
@@ 2.1. Correspondence between places and embeddings
for all $x\in K$, where $|\cdot|_{\ell}$ is the $\ell$-adic absolute value on $\overline{\mathbb{Q}_{\ell}}$ with $|\ell|_{\ell}=\ell^{-1}$. Then, the assignment $\tau\mapsto\lambda_{\tau}$ gives a one-to-one correspondence between embeddings $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$ up to pre-composition with elements in $\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$ and places $\lambda|\ell$ of $K$.
@@
By Lemma 2.1, it suffices to show that the assignment $\sigma\mapsto\iota_{\ell}\circ\sigma$ gives a one-to-one correspondence between embeddings $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$ and embeddings $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$ up to pre-composition with elements in $\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$. Let $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$ be an embedding and let $\tau=\iota_{\ell}\circ\sigma$. Since $\ell$ splits completely in $K$, the completion $\widehat{\tau(K)}$ of $\tau(K)$ in $\overline{\mathbb{Q}_{\ell}}$ is thus the same as $\mathbb{Q}_{\ell}$. In this way, $\tau$ factors as
$\tau:K\xrightarrow{\;\sim\;}\tau(K)\subseteq\widehat{\tau(K)}=\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}},$

### v2

@@ 1. Introduction
Let $A$ be an abelian variety of dimension $g$ over a number field $F$. Let $v$ be a prime of $F$ that is of good reduction for $A$ and let $A_{v}$ denote the reduction of $A$ at $v$. Then, $A$ is said to have ordinary reduction at $v$ if $A_{v}$ is an ordinary abelian variety over $\mathbb{F}_{v}$. Equivalently, if we let $p$ be the rational prime lying below $v$ and let $P_{A,v}(X)$ denote the characteristic polynomial of the Frobenius endomorphism of $A_{v}$, then $A$ has ordinary reduction at $v$ if the coefficient $a_{v,g}(A)$ of $X^{g}$ in $P_{A,v}(X)$ is not divisible by $p$, i.e., $p\nmid a_{v,g}(A)$. We will use the phrases “$A$ has ordinary reduction at vv”, “AA is ordinary at vv”, and “$v$ is an ordinary prime of $A$ over FF” interchangeably. In the case when $A/\mathbb{Q}$ is an elliptic curve, a rational prime $v=p$ is an ordinary prime of $A$ if and only if $p\nmid a_{p,1}(A)$, where $a_{p,1}(A)=p+1-|A_{p}(\mathbb{F}_{p})|$ is the Frobenius trace of $A$ at $p$. When $p\geq 5$, this is also equivalent to that $a_{p,1}(A)\neq 0$ by the Hasse–Weil bound.
@@
A conjecture commonly attributed to Serre [Ser13, #133] predicts that there are infinitely many ordinary primes for $A/F$. Furthermore, it is conjectured that the density of ordinary primes for $A$ over $\overline{F}$ is 1 (see e.g., [Pin98, §7]). The conjecture is known when the dimension of $A$ is 1 or 2 [Ser13, DMOS82, Saw16]. For dimension greater than $2$, the known cases typically assume that $A$ has extra endomorphisms. Our results are also in this direction: we prove that this conjecture holds for certain higher dimensional $\operatorname{GL}_{2}$-type” abelian varieties. In Remark 1.4, we will discuss other known cases of the conjecture in higher dimensions as well as draw comparison to our results.
The main results we have are the following. All the results concern $\operatorname{GL}_{2}$-type” abelian varieties: the first two are for varieties of specific dimensions over $\mathbb{Q}$ and quadratic number fields, respectively, and the last one is for varieties over general number fields with no restriction on the dimension but subject to certain conditions on the splitting behaviors of primes. Throughout, we denote by $\operatorname{End}_{F}(A)$ the ring of $F$-endomorphisms of an abelian variety $A/F$.
@@
    AA is absolutely simple, i.e., $A/\overline{F}$ is simple;
@@
    $A/\overline{F}$ is not of type IV in Albert’s classification.
@@
We make the following remarks on Theorem 1.1, Theorem 1.2, and Theorem 1.3.
  1. (1)
    The “qq” case of Theorem 1.1, Theorem 1.2, and the (P1) case of Theorem 1.3 are in a way immediate consequences of simple analysis of the traces of Frobenius. However, one needs a more detailed analysis of the Galois representations attached to $A$ in order to prove the “2q2q” case of Theorem 1.1 and the (P2) case of Theorem 1.3. In these cases, we first prove a result on the surjectivity of the Galois representations attached to $A$ (see Theorem 3.6) and then use it to deduce the positive density result.
  2. (2)
    In [Fit24], Fité showed that the sets of ordinary primes of certain abelian varieties with extra endomorphisms have positive density. In particular, our results can be viewed as certain generalizations of [Fit24, Theorem 3] from dimension $3$ to certain other dimensions, though the techniques we use are quite different from the ones by Fité.
  3. (3)
    In [Suh20], Suh proved some results on the positive density of the sets of ordinary primes for certain Hilbert modular forms of parallel weight 22.111Hilbert modular forms of parallel weight $2$ conjecturally correspond to certain abelian varieties of $\operatorname{GL}_{2}$-type via some generalization of the Eichler–Shimura theory. One should compare the “2q2q” case of Theorem 1.1 and the (P2) case of Theorem 1.3 with the conditional result [Suh20, Theorem. (d), p.649]. Our proof in these cases ultimately reduces to showing that $\{v:\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}\in\mathbb{Z}\}$ has density 0, where $a_{v}$ is the trace of Frobenius at $v$ of some $2$-dimensional Galois representation attached to $A$. This should follow if one assumes some strong conjectural variant of the Sato–Tate conjecture as considered in [Suh20, Section 3.3]. In our case, we prove this unconditionally by a careful analysis of the image of the Galois representation attached to $A$.
  4. (4)
    The essential parts of our proofs depend only on the properties of the traces of Frobenius and the Galois representations attached to these abelian varieties. We thus formulate those parts separately as Theorem 1.7 and Theorem 1.8. Indeed, one immediate benefit is that one can apply these to obtain similar results on modular forms and Hilbert modular forms (see Theorem 1.6 and Remark 1.9 (2)).
  5. (5)
    Other known cases of the positive density conjecture on the ordinary primes include the case when $A/F$ is a CM abelian variety [Fit24, Remark 12] (i.e., the geometric endomorphism algebra $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ is a CM field of degree $2g$) or when the Mumford–Tate group of $A$ is small in the sense of [Noo95, Theorem 2.2, p. 168] or [Pin98, Theorem 7.1]. More recently, Cantoral Farfán, Li, Mantovan, Pries, and Tang [CFLM+25] established a positive density result for the sets of ordinary primes of certain abelian varieties parametrized by unitary Shimura curves with simple signature.
Remark 1.5.
@@
The family we consider is taken from [TTV91]. Let $p$ be a rational prime with $p\neq 2$ or $5$. Let $\zeta_{p}$ denote the primitive $p$-th root of unity in $\overline{\mathbb{Q}}$, $g(X)\in\mathbb{Z}[X]$ be the minimal polynomial of $(\zeta_{p}+\zeta_{p}^{-1})$, and $f_{t}(X):=Xg(X^{2}-2)+t\in\overline{\mathbb{Q}}[X]$ for any $t\in\overline{\mathbb{Q}}$. The equation
@@
[…]
A comparable notion of “ordinary” also appears in the study of modular forms. Let $f$ be a newform of weight $k\geq 2$ and level $N$, let $a_{n}(f)$ be the $n$-th Fourier coefficient of $f$, and let $K_{f}$ be the coefficient field of $f$. For a rational prime $p\nmid N$ and a prime $\lambda|p$ of $K_{f}$, $f$ is called $\lambda$-ordinary if $\lambda\nmid(a_{p}(f))$, and $f$ is called $p$-ordinary if $f$ is $\lambda$-ordinary for all $\lambda|p$, or equivalently, $\lambda\nmid(a_{p}(f))$ for all primes $\lambda|p$ of $K_{f}$. We will use the phrases “$f$ is $p$-ordinary”, “$f$ is ordinary at pp”, and “$p$ is an ordinary prime of $f$ over $\mathbb{Q}$” interchangeably.
The notion of ordinary modular forms first appeared in a paper by Hida [Hid81]. This paper was followed by a series of works by Hida, where he established the theory of ordinary $p$-adic modular forms, or nowadays known as Hida theory. The ordinary assumption has since played an important role in the theory of ($p$-adic) modular forms, particularly in Iwasawa theory, e.g., in Emerton–Pollack–Weston [EPW06] and Skinner–Urban [SU14] (and many others). The definition of ordinariness that we adopt in this paper is, however, different from (but closely related to) Hida’s definition. Roughly speaking, Hida’s definition of $p$-ordinariness only requires $a_{p}(f)$ to be a $p$-adic unit under a prechosen embedding $K_{f}\hookrightarrow\overline{\mathbb{Q}_{p}}$, while our definition requires $a_{p}(f)$ to be a $p$-adic unit under all embeddings $K_{f}\hookrightarrow\overline{\mathbb{Q}_{p}}$. In particular, the two definitions coincide when $K_{f}=\mathbb{Q}$. The reason for us to adopt a different definition is that a weight $2$ newform $f$ is ordinary at $p$ under our definition if and only if the abelian variety $A_{f}/\mathbb{Q}$ associated to $f$ via the Eichler–Shimura theory is ordinary at $p$ (see Proposition 3.12).
Under this definition, we obtain the following result on ordinary primes of modular forms.
@@
Let $f$ be a newform of weight $2$ and level $N$ and let $K_{f}$ denote the coefficient field of $f$. Suppose that $K_{f}:\mathbb{Q}]=q$ or $2q$ for some prime $q$. Then, the set of ordinary primes of $f$ over $\mathbb{Q}$ has positive density.
@@
As suggested in Remark 1.4 (4), we will first prove the following two results that are formulated independently of abelian varieties or modular forms. Here Theorem 1.7 concerns compatible systems of Galois representations and Theorem 1.8 is merely on sequences of algebraic integers indexed by primes (resembling traces of Frobenius). We will then apply these results to deduce the main theorems for abelian varieties and modular forms. Specifically, Theorem 1.7 deals with the “2q2q” case of Theorem 1.1 and Theorem 1.6, and the (P2) case of Theorem 1.3; Theorem 1.8 deals with the “qq” case of Theorem 1.1, Theorem 1.2, and Theorem 1.6, and the (P1) case of Theorem 1.3.
@@
    (the Ramanujan conjecture/Hasse–Weil bound) the sequence $(a_{v})_{v\in\Sigma_{F}\setminus S}$ satisfies that
@@
    (the Ramanujan conjecture) $|\iota(a_{v})|_{\mathbb{C}}\leq 2\mathrm{N}(v)^{1/2}$ for all $v\in\Sigma_{F}$ with $v\notin S$ and all embeddings $\iota:K\hookrightarrow\mathbb{C}$;
@@
    (non-lacunary) the set $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$ of primes of $F$ has density 0.
@@
We make two remarks on Theorem 1.7 and Theorem 1.8.
  1. (1)
    The surjectivity condition (4) in Theorem 1.7 is intentionally relaxed to avoid the need for assuming that the endomorphisms of the abelian varieties are defined over the base field or that the modular forms have no inner twists.
  2. (2)
    By assuming a condition similar to that in Theorem 1.3 on the existence of some rational prime with specific joint splitting behaviors, we can also apply Theorem 1.7 and Theorem 1.8 to obtain similar positive density results on the ordinary primes of certain Hilbert newforms. The necessary ingredients for the proof in this case are given by [Car86] and [Tay89] on the existence of the associated Galois representations, [Nek12, Appendix B] on the surjectivity of the Galois images, and [Bla06] on the Ramanujan conjecture.
We now end this section by sketching the main steps of the proof of Theorem 1.7 and its application to the (P2) case of Theorem 1.3. The general approach to proving such results can be traced back to the work of Serre, Katz, and Ogus, and has been refined by Pink, Noot, and others, as previously mentioned. Indeed, our proof is based on similar ideas.
  1. Step 1.
    By analyzing the splitting behaviors of primes in (non-Galois) extensions of number fields, construct a positive density set $S_{0}$ of primes $v$ of $F$ that are of degree $1$ and lie above rational primes $p$ that split into two distinct primes in $K$ of the same inertia degree.
  2. Step 2.
    Associate to each prime $v\in S_{0}$ the value $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}$ and use the splitting behavior of $p$ and the Hasse–Weil bound of $a_{v}$ to show that for $v\notin S^{\mathrm{ord}}$, $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}$ belongs to a finite set of integers (with size at most $2^{2g}+1$).
  3. Step 3.
    Use the surjectivity assumption of the residual Galois representations and the Chebotarev density theorem to show that $\{v:\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}=c\}$ for a fixed integer $c$ has density 0, thus concluding that the density of the primes $v\in S_{0}$ with $v\notin S^{\mathrm{ord}}$ is 0.
  4. Step 4.
    Prove a surjectivity result on the Galois representations attached to the relevant abelian varieties in order to apply Theorem 1.7 to the (P2) case of Theorem 1.3.
@@ 2.1. Correspondence between places and embeddings
for all $x\in K$, where $|\cdot|_{\ell}$ is the $\ell$-adic absolute value on $\overline{\mathbb{Q}_{\ell}}$ with $|\ell|_{\ell}=\ell^{-1}$. Then, the assignment $\tau\mapsto\lambda_{\tau}$ gives a one-to-one correspondence between embeddings $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$ up to composition with elements in $\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$ and places $\lambda|\ell$ of $K$.
@@
By Lemma 2.1, it suffices to show that the assignment $\sigma\mapsto\iota_{\ell}\circ\sigma$ gives a one-to-one correspondence between embeddings $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$ and embeddings $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$ up to composition with elements in $\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$. Let $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$ be an embedding and let $\tau=\iota_{\ell}\circ\sigma$. Since $\ell$ splits completely in $K$, the completion $\widehat{\tau(K)}$ of $\tau(K)$ in $\overline{\mathbb{Q}_{\ell}}$ is thus the same as $\mathbb{Q}_{\ell}$. In this way, $\tau$ factors as
$\tau:K\xrightarrow{\;\sim\;}\tau(K)\subseteq\widehat{\tau(K)}$$\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}}$,